OpenStudy (anonymous):

how would i find the 5th roots of -2i? (leave the answers in polar form and the angle in degrees) thank you!

OpenStudy (kinggeorge):

First, you need to change -2i into polar form. Can you show me what you get when it's in polar form?

OpenStudy (anonymous):

Is it 2cis0? I'm not entirely sure.

OpenStudy (kinggeorge):

On a graph, -2i is located there. The angle from the positive real axis, to the negative imaginary axis, is \(3\pi/2\). So the polar form should be \[2\text{ cis} (3\pi/2)\]or\[\Large 2e^{\frac{3i\pi}{2}}\]

OpenStudy (anonymous):

Oh! Okay, I see! I'm going to work it out with the information you have just given me. Thank you!

OpenStudy (kinggeorge):

As a hint, it's a lot easier to take roots of \(2e^{\frac{3i\pi}{2}}\)

OpenStudy (anonymous):

So far, I have 2cis3pi/10 and 2cis7pi/10. Am I on the right track?

OpenStudy (kinggeorge):

Remember that you need to have \[\Large \sqrt[5]{2}\]out in front of those, but otherwise, I think those are correct.